40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10494     Entry time: Thu Sep 11 02:08:32 2014
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Higher transmission powers 

No breakthroughs tonight. 

DRMI didn't want to lock with either the recipe that we used a year ago (elog 9116) or that was used in May (elog 9968).  Being lazy and sleepy, I chickened out and went back to PRFPMI locking. 

Many attempts, I'll highlight 2 here.

(1) I had done the CARM -> sqrtInvTrans transition, and reduced the CARM offset to arm powers of about 7, and lost lock.  I don't remember now if I was trying to transition DARM to AS55, or if I was just prepping (measuring error signal ratio and relative sign).

Zoom_TRXTRYat7_CARMsqrtInvTrans_DARMalsdiff.png

(2) I stopped the carm_cm_up script just before it wanted to do the CARM -> sqrtInvTrans transition, and stayed with CARM and DARM both on ALS.  I got to reasonably high powers, and was measuring the error signal ratios I needed for CARM -> REFL DC and DARM -> AS55.  Things were too noisy to get good coherence for the DARM coefficient, but I thought I was in good shape to transition CARM to REFL DC (which looks like REFL11I, since REFLDC goes to the CM board, and the OUT2 of that board is used to monitor the input to the board. )  Anyhow, I set the offset such that it matched my current CARM offset value, and started the transition, but lost lock about halfway through.  CARM started ringing up here, and I think that's what caused this lockloss.  Could have been the CARM peak, which I wasn't considering / remembering at the time.

ALSonly_TryingTransitionStraightToREFLDC.png

Daytime activity for Thurs:  Lock DRMI, maybe first on 1f signals, but then also on 3f signals.

ELOG V3.1.3-